OpenStudy (anonymous):

Find the measures of the complement and the supplement of an angle with a measure of 20 degree's. A- complement, 150 degrees; supplement, 260 degree's. B- complement, 340 degrees; supplement, 160 degrees C- complement, 70 degrees; supplement, 160 degrees D- complement, 160 degree's; supplement, 70 degrees.

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (anonymous):

complement is 90, so subtract 20 from 90, and supplement is 180, so subtract 20 from 180

OpenStudy (vishweshshrimali5):

Say if I have some angle x

OpenStudy (anonymous):

so 70 and 160?

OpenStudy (vishweshshrimali5):

Now, I want to calculate its y and supplement is z. So, x+y = 90 and x+z = 180

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help me out on that? lol

OpenStudy (vishweshshrimali5):

Now can you find out the answer.

OpenStudy (vishweshshrimali5):

Now, here you have, x = 20. Now, can you find out y and z ?

OpenStudy (anonymous):

yes 70 and 160

OpenStudy (vishweshshrimali5):

@js7474 , did you get it ?

OpenStudy (anonymous):

no not really?

OpenStudy (vishweshshrimali5):

Okay you have x = 20. Now I gave you two equations: For complementary angle (let y) : x + y = 90 For supplementary angle (let z) x+z = 180

OpenStudy (anonymous):

im really bad at this stuff

OpenStudy (vishweshshrimali5):

No problem

OpenStudy (vishweshshrimali5):

Can you tell me the problem you are facing ?

OpenStudy (anonymous):

i have trouble figuring out x and y

OpenStudy (vishweshshrimali5):

Okay here x is the angle given for whose complementary angle and supplementary angle is asked to be calculated. So, here x = 20. Now, y is the required complementary angle and z is the required supplementary angle.

OpenStudy (vishweshshrimali5):

Now can you tell me values of x, y and z ?

OpenStudy (anonymous):

do u still need help b/c I have an easier way to explain it